Price the bonds from the following table with annual coupon payments.

Par Value

Coupon Rate

Years to

Maturity

Yield to

Maturity

Price

​$1,000

10%

10

9%

​$1,000

6%

30

6%

​?

​$5,000

8​%

25

10​%

​?

​$5,000

12​%

20

5%

​?
Answer #1

1)

price=PMT*((1-(1+r)^(-n))/r)+FV/(1+r)^n

=(1000*10%)*((1-(1+9%)^(-10))/9%)+1000/(1+9%)^10

=1064.18

2)

price=(1000*6%)*((1-(1+6%)^(-30))/6%)+1000/(1+6%)^30=1000.00

3)

price=(5000*8%)*((1-(1+10%)^(-25))/10%)+5000/(1+10%)^25=4092.30

4)

price=(5000*12%)*((1-(1+5%)^(-20))/5%)+5000/(1+5%)^20=9361.77

the above is answer..